Im using SPSS to test whether the researchers’ predictions are true and to determine the proportion of cholesterol concentration that is explained by time watching TV. but I dont know what test to run. please help

DATA is below

Case Number

1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
60
61
62
63
64
65
66
67
68
69
70
71
72
73
74
75
76
77
78
79
80
81
82
83
84
85
86
87
88
89
90
92
93
94
95
96
97
98
99
100

Time

168
170
170
164
159
168
165
156
172
170
165
168
171
168
166
167
163
161
168
170
171
165
162
175
169
170
176
179
173
176
170
172
175
172
173
177
170
171
173
170
163
173
169
170
171
174
178
174
176
176
171
176
176
172
173
176
172
171
173
176
171
182
180
183
170
176
182
174
178
176
169
172
176
175
170
177
180
176
171
183
177
169
167
171
170
174
170
170
173
173
177
168
179
163
179
168
170
171
171

Cholestrol

4.60
4.80
5.39
5.16
5.09
5.70
5.25
4.89
4.90
4.68
4.77
4.65
5.61
4.81
5.64
6.16
5.35
4.65
5.79
4.80
5.05
5.66
4.63
5.01
5.56
6.41
5.03
5.35
5.09
5.99
4.72
4.70
5.01
5.25
5.83
6.06
6.10
5.27
5.54
5.84
5.01
6.04
4.69
5.43
5.71
6.41
5.26
5.61
6.11
5.25
5.53
6.55
5.25
5.55
5.50
5.23
5.12
6.27
5.31
5.10
5.12
5.91
5.91
6.33
4.99
5.10
5.56
5.27
5.60
4.87
5.73
4.58
5.22
5.13
5.68
6.21
6.21
5.99
4.57
5.66
5.23
5.82
6.14
4.80
5.23
5.91
6.21
4.65
5.56
5.16
5.65
5.30
6.00
5.48
6.43
4.36
5.17
5.52
6.10
Answer #1

SLOUTION :

EXCEL SCREEN SHOT:

Page Layout Formulas Data Home Insert Review View Cut 11A A Calibri Wrap Text General Copy Paste ▼ Or▼ 스 ▼ % , 넒4°: Condition

y^ = -0.9365 + 0.0370*x

we want to test

Ho : B = 0

Ha :61> 0

p-value = 0.0003/2 = 0.0001 { as this is one-tailed test}

since p-value < alpha

we reject the null hypothesis

we conclude that there is a positive relationship

the proportion of cholesterol concentration that is explained by time watching = R^2 = 0.1292

= 12.92%